CNNBC recently reported that the mean annual cost of auto insurance is 996 dollars. Assume the standard deviation is 275 dollars. You take a simple random sample of 68 auto insurance policies.

Find the probability that a single randomly selected value is less than 975 dollars.
P(X < 975) =

Find the probability that a sample of size n=68n=68 is randomly selected with a mean less than 975 dollars.
P(M < 975) =

Enter your answers as numbers accurate to 4 decimal places.
Answer #1

P = 996 , q = 276 , n = 6 8 P(x <975 ) = P(Z < 975 - 996) (a) 275 = p (75 -0.08) -0.4681 HR (0) P(x2975) = P(28 975-999) [for